This question has been answered incorrectly twice. Please do not do if you do not know how to do. Thank you.

Improvised explosive devices (IEDs) are responsible for many deaths in times of strife and war. Unmanned ground vehicles (robots) can be used to disarm the IEDs and perform other tasks as well. If the robots cost $155,000 each and the military arms unit signs a contract to purchase 4,000 of them now and another 7,500 one year from now, what is the equivalent annual cost of the contract over a 8-year period at 8% per year interest? The equivalent cost of the contract is determined to be $
Answer #1

Let the equivalent annual cost be X

Hence, NPV of annual costs for 8 years = X + X/(1+r) + X/(1+r)2 +.....+ X/(1+r)7

= X + X[1 - (1+r)-6]/r

Given r = 0.08

=> NPV = 5.62X

Now, this will be equal to the NPV of the payment made this year and payment to be made next year

=> 5.62X = 155000*4000 + 155000*7500/(1.08)

=> X = 301,741,175.54

Hence, equivalent annual cost = $301,741,175.54